OpenStudy (anonymous):

Please help!!!! I'm confused on how to solve this problem. r=radical sign Evaluate: 5r8 r8/3r8^5 In words... The fifth root of 8 and square root of 8 divided by the third rot of 8 raised to the 5th power

OpenStudy (anonymous):

@Phi @Hero @across @amistre64

OpenStudy (aum):

\[ \Large \frac{\sqrt[5]{8}*\sqrt{8}}{(\sqrt[3]{8})^5} = \frac{8^{1/5}*8^{1/2}}{(8^{1/3})^5} = \frac{8^{1/5+1/2}}{8^{5/3}} = 8^{1/5+1/2-5/3} = ? \]

OpenStudy (anonymous):

Thx but I really need an answer, the question is multiple choice so here are the given answers A: 8 -3/10 B: 8 -41/30 C: 8 -29/30 D: 8 71/30 They are fractions.

OpenStudy (aum):

Not allowed to give out just answers. You should be able to add the fractions: \(\Large \frac 15 + \frac 12 - \frac 53 = ?\).

OpenStudy (anonymous):

Oh I simply thought you didn't know the answer to my problem...anyhow I got it its C because once added they are -29/30. thank you so much

OpenStudy (anonymous):

Welcome to OpenStudy :)